OpenStudy (theoreo):

Can somebody show me how to solve cos(x)+cos(2x) = -1 I know the answer is (pi/2) but when I tried to solve it, I got (pi/3), and I'm not sure how to do it right.

OpenStudy (perl):

you can expand cos(2x)

OpenStudy (perl):

cos(2x) = cosx * cos x - sin x * sin x = cos^2(x) - sin^2(x) substitute 1-cos^2(x) for sin^2(x) = cos^2(x) - (1 - cos^2(x)

OpenStudy (theoreo):

Ah, ok. Then solve it like a quadratic?

OpenStudy (perl):

correct

OpenStudy (theoreo):

Ok great. Thanks a lot.

OpenStudy (dan815):

yes good work men

OpenStudy (perl):

dan :D

OpenStudy (anonymous):

perl :D

OpenStudy (dan815):

baby:d

OpenStudy (perl):

hey batman

OpenStudy (anonymous):

Congrats on 99

OpenStudy (perl):

thanks :)